# Understanding Emotion in Discourse: Recognition Insights and Linguistic Patterns for Generation

### v1

@@ 6 Limitations and Future Directions
While our findings reveal consistent and interpretable patterns in how models encode emotion, several limitations should be acknowledged and motivate future extensions.
@@
Although the positional effects we report are statistically reliable, their effect sizes remain modest (Cramér’s $V=0.062$). Emotional expression is inherently variable across speakers and contexts, and our observed patterns should be interpreted as probabilistic tendencies rather than deterministic linguistic laws. Future work should test whether these discourse-level patterns generalize across more diverse corpora such as MELD (poria2018meld), CMU-MOSI (zadeh2016mosi), and CMU-MOSEI (zadeh2018multimodal).
@@
Our analysis is limited to text-based emotion recognition, which captures linguistic and discourse-level cues but omits prosodic and visual signals. Given that emotions like Excited are distinguished primarily through acoustic intensity, extending this framework to multimodal data would enable precise quantification of where and how much information from prosody or facial dynamics contributes beyond text.
Dataset bias and generalizability.
Because IEMOCAP contains acted English dialogues, its label distributions and emotional dynamics may not reflect spontaneous or cross-linguistic behavior. The emotion-specific context patterns we observed (e.g., Sad requiring extended context while Angry saturates quickly) might thus be dataset-specific artifacts. Cross-dataset and multilingual replication will be essential to determine whether these represent general computational principles or culturally contingent effects.
Discourse marker inventory.
Our analysis relies on a predefined 20-marker inventory drawn from established taxonomies. While these markers are well-attested in the linguistics literature, this closed set may miss emotion-specific expressions or informal markers prevalent in conversational speech. Data-driven marker discovery could complement our theory-driven approach.
@@ 7 Conclusion
We presented a systematic analysis of emotion recognition in conversation, addressing two gaps in the literature: understanding which architectural choices matter for recognition, and identifying linguistic patterns that could inform generation.
For recognition, our experiments reveal three key findings. First, conversational context is paramount: incorporating preceding dialogue yields substantial improvements, with our past-only approach achieving state-of-the-art results (82.69% for 4-way, 67.07% for 6-way) among text-only methods. Second, hierarchical sentence encoding helps at utterance-level but provides no benefit once conversational context is available, suggesting that turn-level dynamics subsume intra-utterance structure. Third, external affective lexicons (SenticNet) provide no improvement, indicating that pre-trained encoders already capture sufficient emotional semantics.
For linguistic analysis, our examination of 5,286 discourse marker occurrences reveals emotion-specific positioning patterns. Sad utterances show significantly reduced left-periphery marker usage (21.9% vs. 28–32% for other emotions), consistent with diminished turn-management behavior associated with low-arousal withdrawal. This connects to our finding that Sad benefits most from context (+22%p): lacking explicit pragmatic markers, sadness must be inferred from conversational trajectory.
Our confusion analysis further questions the validity of 6-way emotion taxonomies for text-only classification. The asymmetric Happy–Excited confusion (20.6% vs. 1.3%) and unexpectedly high cross-valence errors suggest that categorical boundaries may not align with linguistic distinguishability.
These findings provide actionable insights for ERC system design: simple flat encoders with moderate context (K=30K=305050 preceding turns) capture most predictive information, position-weighted pooling helps only without context, and 4-way classification may be more appropriate for text-only applications. For emotion-conditioned generation, our discourse marker findings suggest concrete production guidelines, such as reducing turn-initial markers for sad utterances.
@@ Appendix A Discourse Marker Inventory
Our discourse marker analysis uses markers drawn from established taxonomies in discourse and pragmatics research. We compiled a search inventory from schiffrin1987discourse, fraser1999discourse, traugott2010intersubjectivity, verhagen2005constructions, aijmer2013understanding, beeching2014discourse, and biber1989styles. Table 6 lists the 20 markers that were empirically found in IEMOCAP, along with their frequencies and theoretical sources.
@@
and Elaborative 2,372 schiffrin1987discourse; fraser1999discourse
so Inferential 1,968 schiffrin1987discourse; fraser1999discourse; beeching2014discourse
like Pragmatic particle 1,210 aijmer2013understanding
but Contrastive 760 schiffrin1987discourse; fraser1999discourse; beeching2014discourse
well Turn-management 727 schiffrin1987discourse; beeching2014discourse
oh Turn-management 564 schiffrin1987discourse; beeching2014discourse
you know Intersubjective 393 schiffrin1987discourse; verhagen2005constructions
i mean Intersubjective 240 schiffrin1987discourse; verhagen2005constructions
maybe Epistemic (doubt) 195 traugott2010intersubjectivity; biber1989styles
though Contrastive 162 fraser1999discourse; beeching2014discourse
i think Epistemic (stance) 131 traugott2010intersubjectivity
probably Epistemic (doubt) 83 traugott2010intersubjectivity; biber1989styles
i guess Epistemic (stance) 77 traugott2010intersubjectivity
yet Contrastive 29 fraser1999discourse
also Elaborative 18 fraser1999discourse
i believe Epistemic (stance) 10 traugott2010intersubjectivity
however Contrastive 6 fraser1999discourse
although Contrastive 5 fraser1999discourse
unfortunately Attitudinal 4 biber1989styles
therefore Inferential 1 fraser1999discourse

### v2

@@ 5 Limitations and Future Directions
While our analyses reveal consistent and interpretable patterns in how causal ERC models leverage context and discourse cues, several limitations should be acknowledged and motivate future extensions.
@@
The discourse-marker effects we report are statistically reliable but small in magnitude (e.g., Cramér’s $V=0.062$). Emotional expression is highly variable across speakers, interactional goals, and situations; accordingly, our positional patterns should be interpreted as probabilistic tendencies rather than deterministic linguistic laws. Future work should quantify robustness under domain shift and across interaction types (e.g., multi-party chat in MELD (Poria et al., 2018), written dialogue in DailyDialog, and other conversational corpora), and should report effect sizes alongside significance to avoid overinterpretation.
@@
Our experiments are restricted to text-based ERC. This captures lexical and discourse-level cues but omits prosodic and visual signals that are often crucial for separating closely related categories (e.g., excited vs. happy). Extending the same ablation and fixed-context protocol to multimodal encoders would allow precise attribution of what additional information is contributed by acoustics and facial dynamics beyond text.
Dataset bias, label noise, and generalizability.
IEMOCAP consists of acted English dyadic dialogues, which may differ from spontaneous interaction in turn-taking style, emotional intensity, and marker usage. In addition, categorical emotion labels in conversational corpora are inherently noisy and sometimes underspecified for text-only interpretation. The emotion-specific context patterns we observe (e.g., larger gains for Sad than Angry) may therefore partially reflect dataset- and annotation-specific properties. Cross-dataset replication and multilingual evaluation are essential to determine whether these are general computational regularities or corpus-contingent effects.
Operationalization choices in context and discourse analysis.
Our context analysis uses turn-based windows and a fixed definition of “past-only” context, and our discourse analysis relies on normalized token positions with predefined periphery thresholds (LP <0.15<0.15, RP >0.85>0.85). Alternative operationalizations (e.g., sentence- or time-based windows, speaker-conditioned windows, syntactic peripheries, or discourse-unit segmentation) may yield different quantitative estimates. Future work should stress-test these choices and assess whether conclusions hold under reasonable variants.
Closed discourse-marker inventory.
We analyze a predefined inventory of 20 markers drawn from established taxonomies. While theory-driven, this closed set may miss informal and multiword pragmatic expressions prevalent in conversational speech. A promising direction is to complement this with data-driven discovery (e.g., collocation-based mining, pragmatic phrase induction, or supervised taggers) and to examine whether automatically discovered marker families show stronger or more generalizable emotion associations.
@@ 6 Conclusion
We presented a systematic analysis of emotion recognition in conversation (ERC) that targets two gaps in the literature: (i) clarifying which modeling choices materially affect recognition under a strictly causal (past-only) setting, and (ii) identifying interpretable discourse patterns that can inform future work on emotion-conditioned generation.
For recognition, our experiments yield three primary takeaways. First, conversational context is the dominant driver of performance: most of the attainable gain is achieved with a relatively short window of preceding turns, after which returns diminish. Second, intra-utterance hierarchical sentence encoding can be beneficial when no conversational context is available, but its advantage largely disappears once turn-level context is introduced, suggesting that inter-turn dynamics can substitute for fine-grained within-utterance structure in causal ERC. Third, integrating an external affective lexicon (SenticNet) provides no improvement in our setting, consistent with the view that modern pretrained encoders already capture much of the affective semantics needed for classification.
For linguistic analysis, our corpus study of 5,286 discourse-marker occurrences reveals small but reliable emotion-specific positional tendencies. In particular, Sad utterances exhibit reduced left-periphery marker usage relative to other emotions, which we interpret as diminished overt turn-management signaling. This observation complements the recognition-side finding that Sad benefits most from added context: when explicit pragmatic cues are weaker, conversational history becomes more important for disambiguation.
Finally, our confusion analysis highlights that, in text-only ERC, some fine-grained distinctions in the 6-way taxonomy are weakly identified without prosodic information (e.g., asymmetric Happy–Excited confusions). Taken together, these results suggest practical guidance for causal ERC system design: simple flat encoders with moderate past context can capture most predictive information; more elaborate utterance structure offers limited marginal benefit once context is available; and careful attention to discourse phenomena offers a principled path toward interpretability and generation-relevant insights.
Acknowledgements
This work was supported by the UC Noyce Initiative. We thank the anonymous reviewers for their helpful comments. Portions of the writing and visualizations were developed with assistance from Claude (Anthropic). The authors remain fully responsible for all content.
@@ Appendix A Discourse Marker Inventory
Our discourse marker analysis uses markers drawn from established taxonomies in discourse and pragmatics research. We compiled a search inventory from Schiffrin (1987), Fraser (1999), Traugott (2010), Verhagen (2005), Aijmer (2013), Beeching and Detges (2014), and Biber and Finegan (1989). Table 5 lists the 20 markers that were empirically found in IEMOCAP, along with their frequencies and theoretical sources.
@@
and Elaborative 2,372 Schiffrin (1987); Fraser (1999)
so Inferential 1,968 Schiffrin (1987); Fraser (1999); Beeching and Detges (2014)
like Pragmatic particle 1,210 Aijmer (2013)
but Contrastive 760 Schiffrin (1987); Fraser (1999); Beeching and Detges (2014)
well Turn-management 727 Schiffrin (1987); Beeching and Detges (2014)
oh Turn-management 564 Schiffrin (1987); Beeching and Detges (2014)
you know Intersubjective 393 Schiffrin (1987); Verhagen (2005)
i mean Intersubjective 240 Schiffrin (1987); Verhagen (2005)
maybe Epistemic (doubt) 195 Traugott (2010); Biber and Finegan (1989)
though Contrastive 162 Fraser (1999); Beeching and Detges (2014)
i think Epistemic (stance) 131 Traugott (2010)
probably Epistemic (doubt) 83 Traugott (2010); Biber and Finegan (1989)
i guess Epistemic (stance) 77 Traugott (2010)
yet Contrastive 29 Fraser (1999)
also Elaborative 18 Fraser (1999)
i believe Epistemic (stance) 10 Traugott (2010)
however Contrastive 6 Fraser (1999)
although Contrastive 5 Fraser (1999)
unfortunately Attitudinal 4 Biber and Finegan (1989)
therefore Inferential 1 Fraser (1999)